OpenStudy (anonymous):

I have a right triange with sides of x and y. I know y is a constant (500) and that dtheta/dt (where theta is the angle opposite from side x) is also constant (8pi rad/s). I have already found dx/dt for a specified x, but the next problem asks me to relate dx/dt to x. How do I do this?

OpenStudy (anonymous):

if i am picturing it right \[\tan(\theta)=\frac{x}{y}=\frac{x}{500}\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Sorry for the formatting; I hadn't figured out the interface when I posted...

OpenStudy (anonymous):

i guess i am confused if you found \(\frac{dx}{dt}\) for a specific \(x\) why can't you find it for a general \(x\)?

OpenStudy (anonymous):

I may have just got it... thinking...

OpenStudy (anonymous):

i think you are being asked to put \(x\) where you put the number you used for \(x\) and write that i think

OpenStudy (anonymous):

Got it. \[\frac{dx}{dt}=\frac{4\pi d^{2}}{250}+100\] When I differentiate the original problem, I get an answer in terms of \(\theta\) and \(\frac{d\theta}{dt}\). \(\frac{d\theta}{dt}\) is known and theta can be found in terms of \(x\) because of pythagorean theorem stuff. I found \(\theta\) in terms of the hypotenuse and then the hypotentus in terms of \(x\).

OpenStudy (anonymous):

good work but why is that 100 there?

OpenStudy (anonymous):

maybe i did it wrong' i have \[\tan(\theta)=\frac{x}{500}\] so \[\sec^2(\theta)\theta '=\frac{x'}{500}\]

OpenStudy (anonymous):

Oops... it is actually 100. \[\frac{dx}{dt}=\frac{4\pi h^{2}}{250}\] and \[h^{2}=x^{2}+y{2}\] so \(h^{2}\) is replaced with \(x^{2}+250000\) and then the fraction is split to cancel the 250 from the denominator

OpenStudy (anonymous):

oops... it is actually 1000 /me facepalms

OpenStudy (anonymous):

okay